RUBIDIUM ATOMS ON HELIUM DROPLETS: ANALYSIS OF AN EXOTIC RYDBERG COMPLEX

\underline{FLORIAN LACKNER}, GÜNTER KROIS, MARKUS KOCH, and WOLFGANG E. ERNST, Institute of Experimental Physics, Graz University of Technology, Petersgasse 16, A-8010 Graz, Austria.

Rubidium atoms on the surface of superfluid helium droplets have been excited into Rydberg states. The excitation spectrum of the Rb-He_{N} system has been recorded from the 5^{2}D state manifold up to the ionization threshold^{a} by resonant three-photon-ionization time-of-flight spectroscopy^{b}. The observation of droplet size dependent shifts of excited states with respect to bare atom states is explained by a decreased quantum defect and a lowered ionization threshold. Within the scope of a Rydberg model^{c} we demonstrate that quantum defects and ionization thresholds are constant for each specific Rydberg series, which confirms the Rydberg character of excited Rubidium states on helium droplets. A set of six Rydberg series could be identified. Individual Rydberg states are observed with effective principle quantum numbers up to n^{∗} ≈ 19 and l ≤ 3, for which the expectation value of the electron orbital radius is about ten times larger than the droplet radius.

aM. Theisen, F. Lackner, G. Krois, and W.E. Ernst, J. Phys. Chem. Lett., 2, 2778-2782 (2011)

bF. Lackner, G. Krois, M. Theisen, M. Koch, and W.E. Ernst, Phys. Chem. Chem. Phys., 13, 18781-18788 (2011)

cJ.E. Murphy, J.M. Berg, A.J. Merer, N.A. Harris, and R.W. Field, Phys. Rev. Lett. 65, 1861 (1990)